OpenStudy (anonymous):

a square plot of land is to be seeded with grass seed. if the plot is 22 meters by 22 meters and a 1 kilogram bag of seed is needed for 85 square meters of land, how many bags of seed are needed?

OpenStudy (anonymous):

i came up with two bags but it seemed to simple

OpenStudy (anonymous):

Six bags

OpenStudy (anonymous):

22*22=484 square meters 484/85= number of bags needed

OpenStudy (anonymous):

multiply the dimensions out (22 and 22) then divide the result by the amount of coverage per bag (85)

OpenStudy (anonymous):

1 sq km needs 1/85th of a bag of seeds. Therefore 22x22 needs 484/85 bags

OpenStudy (anonymous):

approx. 5.694 bags

OpenStudy (anonymous):

22*22=484 Sq meters of land 484/85 5.69, if you can only get 1 kilo bag you would need six (would have seeds leftover)

OpenStudy (anonymous):

Probably 6 bags if one cannot slice bags into fractions of a bag.